OpenStudy (anonymous):

Find the equation of the line containing the given point and slope. (-4, 1) m=2/5

OpenStudy (anonymous):

Any particular form of the equation? I'd use point-slope form: \[(y-y_{1}=m(x-x_{1})\] where the given point is \[(x_{1},y_{1})\] and m is the given slope.

OpenStudy (anonymous):

Sorry...remove the open parenthesis at the front of the point-slope form.

OpenStudy (anonymous):

I you need it in slope-intercept, solve the above answer for y (after plugging in the values).

OpenStudy (anonymous):

yah I tried that and I must of distributed something wrong

OpenStudy (anonymous):

can you post your work? I can look to see where you might have goofed.

OpenStudy (anonymous):

When I distributed 2/5 to x-(-4) i got 2/5x -8/5 which is where i probably got something wrong

OpenStudy (anonymous):

Ahh...x-(-4) = x+4. so you should get 2/5x+8/5. Not switching the sign when subtracting a negative is an all too common mistake.

OpenStudy (anonymous):

ok, its still showing wrong, all they want is they equation

OpenStudy (anonymous):

I get y=(2/5)x + (13/5)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw :)